# Effects of the entropy source on Monte Carlo simulations

### v1

@@ 2 Results
We have demonstrated differences in the sampling itself, as well as the potential of QRNGs to yield performance advantages in terms of improved accuracy or reduced sample size in examplary MC simulations. Reductions in sample size without loss of accuracy should benefit every stochastic simulation (e.g., Black-Scholes, stochastic molecular dynamics), but we expect greatest benefits for problems where random numbers are essential, but do not dominate the overall computation time. Additionally, given the observations of an increased homogeneity of samples, we expect non-linear problems to enhance the differences between the random number generators.
Stochastic Schrödinger equation
Therefore, we turn to a non-linear example with an application in high-accuracy measurements, critical for various fields including physical measurements, biomedical sensing, or precision manufacturing: we investigate a model based on the stochastic Schrödinger equation (SSE) for quantum-enhanced measurements.
As an example we consider the measurement of the relative change $\frac{\delta L}{L}$ in the length of an optical cavity in which two groups of $N/2$ excited atoms are deposited. It is possible to store the two groups in a way to suppress the emission of light into the cavity mode (light frequency), keeping the atoms excited. Once the preparation is done, the length of the cavity is changed, thus changing the coupling constants $G_{i}$ of group $i=1,2$ and resulting in the emission of light.
Braun and Martin have shown in [8] that counting the number of emitted photons allows for the best-possible precision of the measurement of the change of the cavity length $x\propto\frac{\delta L}{L}$, following the Heisenberg limit $\sim N^{-1}$. To double the precision of the measurement, we thus would have to double the number of atoms in the cavity. In contrast, the standard quantum limit scaling $\sim N^{-1/2}$ would require us to quadruple the number of atoms for the same precision.
For small systems (fewer than $\approx 22$ atoms) the state $\psi_{t}$ of the system can be determined using exact diagonalisation (ED) applied to the Lindblad-matrices of the associated master equation [8, eq. (19)]. The procedure becomes unfeasible at system sizes beyond $\sim$ 30 atoms due to the exponentially growing memory requirements, leaving only the stochastic Schrödinger equation as a feasible way to explore the behaviour for large systems (c.f SM sec. A.5).
In this work, we approximate the photon number as a function of dimensionless time $gt$ by numerically solving stochastic Schrödinger equation using a Runge-Kutta-type integrator scheme with strong order-$1$ convergence [13, Ch. 12.3]. The approximations for $N=12$ particles and $M_{steps}=20000$ time steps in the interval $[0,20]$ and varying number of repetitions (number of $\psi_{t}$-trajectories simulated), $N_{rep}$, are compared between QRNG and Mersenne Twister PRNG input, and benchmarked against the analytical solution obtained using exact diagonalisation (ED). The validity of a short-time analytic approximation of the uncertainty used in [8] limited the time of interest there to $gt\leq 0.2$. In practice the measurements must be performed outside of the cavity and such short-time measurements can be challenging and thus longer observation times are preferred.
Figure 4: Simulation of the average photon number for a system with $N=12$ atoms using the stochastic Schrödinger equation and $N_{rep}=50$ trajectories. The trajectories are plotted in A) for PRNG input and B) for QRNG input. The respective standardised error of the means of the trajectories referenced to the analytical solution is shown in C. D) shows histograms at the important long-time limit. E) Standardised deviation of the photon number in the long-time limit obtained with $N_{rep}=50$ for a system of $N=20$ atoms.
Here, fig. 4 illustrates the distribution of $N_{rep}=50$ trajectories of the expected number of photons $\langle n_{ph}\rangle(t)/x^{2}$ as a 2D histogram for simulations performed using the Mersenne Twister PRNG (A) and a QRNG (B). The deviation of the means from the reference solution (obtained using the ED solution), normalised by the standard deviation ($\sigma[\langle n_{ph}\rangle]$) - $\frac{|\mathbb{E}[\langle n_{ph}\rangle]-\langle n_{ph}\rangle_{ED}|}{\sigma[\langle n_{ph}\rangle]}$ taken over all trajectories at each time for the system size $N=12$ is shown in fig. 4C.
The initial conditions are identical across all simulations: each trajectory starts out with a fully excited system at $gt=0$. As the system evolves, the excitation decays by emission of light, resulting in the observable peak in photon numbers. The maximum number of photons is observed at $T_{peak}=gt\approx\frac{7}{10}$ before the system decays into a metastable state in the long-time limit $gt>5$. Interestingly, the standardised deviation of the means from the reference solution is smaller for both the peak time, as well as in the long-time limit for the QRNG input (see fig. 4 C). The latter limit is of particular interest, since, in reality, measurements must be performed outside of the cavity and the longer the time scale for measurements is, the easier these are to carry out. Additionally, according to photon statistics, we expect the distribution of the trajectories at a given time to resemble the Poisson distribution with a variance $\sim\langle n_{ph}\rangle$. Thus, in the long-time limit, we expect to sample the distribution better than at the early times due to a much lower $\sim\langle n_{ph}\rangle$, allowing us to not only assess a single numerical measure (median) but also the shape of the obtained distribution. We must further remark that an evaluation using the standardised deviation is applied, in general, to normally distributed data (i.e., to perform a t-test) and should be carefully scrutinised here, since photon numbers are expected to follow a Poisson distribution. This resemblance to a Poisson distribution is enhanced in the long-time limit ($gt=13$) if outliers, according to the inter-quartile-range (IQR) measure, are truncated from the dataset. We do not include outlier analysis in this work due to the complexity of the subject, but remark that the simple classification according to the IQR is challenging as it can yield acceptance ranges including negative values, which are physically incorrect.
One observes that the QRNG and PRNG data capture the true solution within the range of their values. However, while the median of the QRNG data coincides with the true solution, the median of the PRNG trajectories differs from the true solution by a bin width (c.f. fig. 4 D). Here, the width of the bins of the QRNG data is significantly larger than that of the bins of the PRNG data, which is due to the presence of the singular outlier trajectory in QRNG data. Excluding the singular bin results in a distribution with a smaller variance for QRNG than PRNG trajectories, but, due to its definition, does not modify the median.
Next, it is worth noting that the SSE is supposed to simulate large systems. Similar to deterministic Runge-Kutta schemes, the approximation error is dependent on the time step size. Computing the approximation error w.r.t. the ED solution for system sizes $N=2\dots 16$ and varying time step sizes, one can observe that the step size at which the lowest relative error is achieved scales as $\sim N^{-2}$, which results in an optimal step size $g\Delta t=10^{-3}$ for systems $N\leq 14$ and $g\Delta t=10^{-4}$ for $N\sim 20$  [^4] Adapted from unpublished, prior work by one of the authors.. This is particularly noticeable in the approximation error of the PRNG dataset in fig. 4E, where we provide the long-time limit behaviour of the standardised deviation for $N=20$. It is immediately apparent that using either $\Delta t$, the QRNG error is smaller than the PRNG error. Surprisingly, the PRNG error at $g\Delta t=10^{-4}$ is only marginally better than the QRNG error at $g\Delta t=10^{-3}$. It is important to note that the simulation cost (time) scales linearly with the number of time steps. Hence, the tentative implication of this observation is that the same quality of the approximation with the PRNG can be achieved using the QRNG but at $1/10$-th the cost, using fewer but larger time steps.
Figure 5: Simulation of the average photon number for a system with $N=12$ atoms using the stochastic Schrödinger equation and $N_{rep}=150$ trajectories. The trajectories are plotted in A) for PRNG input and B) for QRNG input. C) The histograms at the long-time limit is illustrated. D) The respective standardised error of the means of the trajectories referenced to the analytical solution is shown.
Increasing the number of trajectories $N_{rep}$ to $150$ reduces the late-time deviation of the PRNG mean to the same order of magnitude as that of the QRNG dataset (c.f. fig. 5 D). Here, in the important long-time limit, where both QRNG and PRNG yield essentially equivalent results, though the QRNG median at $gt=13$ yields a better approximation of the ED solution (c.f. fig. 5 C). Significantly, tripling the number of trajectories to $N_{rep}=150$ has reduced the standardised error of the PRNG dataset to the magnitude of the error of the QRNG dataset already achieved at $N_{rep}=50$.
In summary, considering the temporal evolution of the scaled deviation and the trajectory distributions at $gt=13$, we may conclude that using a QRNG provides a better result than the MT PRNG. The two generators become essentially equivalent if the number of trajectories simulated grows beyond $N_{rep}\approx 150$ for the given system. The long time limit deviation of the PRNG is then reduced to $\approx 10^{-1}$, i.e., the same value obtained with the QRNG input at only $50$ trajectories. This implies that equivalent results can be achieved using QRNGs at $1/3$ the cost. Additionally, when inspecting the behaviour for larger systems, we conclude that the use of a QRNG makes the system less susceptible to sub-optimal choice of the time step size, yielding equivalent results at potentially $1/10$ of the cost.
@@ 3 Conclusions
Likewise, using Buffon’s needle experiment, we demonstrated that the QRNG results pass the t-test (hypothesis: mean = $\pi$, data normally distributed) for sample sizes up to 4.54×4.54$\times$4.54 × larger than those achieved by the parallel PRNG, thus resulting in a 2$\sim$ 2$\times$ 2 × better approximation of $\pi$. If the limit is relaxed to $1/\sqrt{N}$, QRNG permits up to 3.16×3.16$\times$3.16 × larger samples, resulting in an up to 2.9×2.9$\times$2.9 × better approximation. In the process, we have demonstrated the capability of this method to distinguish the effects of points-per-sample $M$ and sample number $N$, which is not commonly found in the literature. We have also shown that the use of a QRNG results in better approximations (lower approximation errors) for sub-optimal parameter choices ($N\gg M$), enabling large reductions (8$\sim$ 8$\times$ 8 ×) of the number of samples without loss of accuracy.
Additionally, based on a uniformity analysis, we assessed differences in the random sampling underpinning the MC simulations: our findings suggest that the QRNG, especially at small sample sizes, offers a better dispersion of samples than the PRNG indicating a tendency of the QRNG towards more uniformly distributed samples, closely matching theoretical expectation.
We further investigated one applied example using the stochastic Schrödinger equation, where a better accuracy of the approximations with the QRNG was revealed by assessing the standardised error from the analytical solution in the relevant long-time limit. The approximation error is smaller when using the QRNG as opposed to the PRNG. It was found that the errors of the PRNG data from the true solution are in the same order of magnitude of the errors of the QRNG data even if the number of trajectories of the latter is reduced by a factor of 3. This implies that the simulation can be run at $\frac{1}{3}$-rd of the costs when using a QRNG. When simulating larger systems, the time-step size must be reduced, resulting in an increase in the number of time steps. It becomes evident that the acceptable approximation error of the PRNG is already achieved by the QRNG data with 1/10-th of the number of time steps. Significantly, the simulation time, and thus cost, scales linearly with the number of time steps.
Moving beyond the investigated use cases with known solutions, we envisage, and encourage, further testing and applications of QRNGs in stochastic modelling, ranging from Bayesian inference, stochastic differential equations, optimisation, and Monte Carlo simulations to leverage the revealed advantages in all affected fields. We plan to pursue a wider comparison of QRNGs, classical true random number generators (TRNGs) and a variety of PRNGs in the aforementioned fields using the methodology utilised in this paper.
@@ A.4 Statistical tests
A.5 Stochastic Schrödinger equation
Braun and Martin have shown in [9] that counting the number of emitted photons allows for the best-possible precision of the measurement of the change of the cavity length $x\propto\frac{\delta L}{L}$, following the Heisenberg limit $\sim N^{-1}$. A numerical verification of the statements requires the calculation of
[…]
Here $n_{ph}$ is the number of photons and $g=|G_{1}-G_{2}|/(2x)$, where $G_{1,2}=1.0\pm x$ are the coupling constants for the two groups of atoms using the photon escape rate $\kappa=5g$. The coupling parameter $x=0.1\propto\frac{\Delta L}{L}$ encodes the relative change $\Delta L$ in the length $L$ of the optical cavity. $J_{\pm}$ are the the generalised spin raising/lowering matrices that represent excitation/relaxation of a spin by way of absorption/emission of a photon.
Herein, the expectation values are computed at a given time $t$ by computing $\langle J_{+}J_{-}\rangle_{\psi_{t}}=\langle\psi(t)|J_{+}J_{-}|\psi(t)\rangle$ with the state $|\psi_{t}\rangle=|\psi(t)\rangle$ obtained using the stochastic Schrödinger equation:
$d|\psi_{t}\rangle$ $=F_{1}(|\psi_{t}\rangle)dt+F_{2}(|\psi_{t}\rangle)dW_{t}$ (11)
$F_{1}(|\psi_{t}\rangle)$ $=\frac{g^{2}}{\kappa}\left(2\langle J_{i}\rangle_{\psi_{t}}J_{-}-J_{+}J_{-}-\langle J_{-}\rangle^{2}_{\psi_{t}}\right)|\psi_{t}\rangle$ (12)
$F_{2}(|\psi_{t}\rangle)$ $=\sqrt{2\frac{g^{2}}{\kappa}}\left(J_{-}-\langle J_{-}\rangle_{\psi_{t}}\right)|\psi_{t}\rangle$ (13)
where $dW_{t}=dW(t)$ is a Wiener increment with zero mean and variance $dt$.
@@ B.1 On the number $\pi$ approximations
Similarly, the normalised deviation displayed in fig. 9 demonstrates that, up to $10^{5}$ points per sample (pps), this methods yields better results. The approximation error of method A displays no clear dependence on $M$ and is hence assumed independent of $M$. In contrast, Buffon’s method yields larger approximation errors for $M\leq 5\cdot 10^{5}$ but displays the expected $\frac{1}{\sqrt{M}}$ convergence, resulting in a better approximation at $M=10^{6}$ pps and beyond. In conjunction with the analysis presented below, we conclude that method B allows for a clear distinction of the effects of the sample size $N$ and the points per sample $M$, whereas for method A such a distinction is not easily made (fig. 9).
@@ B.2 Uniformity analysis - Additional results
W test T test AD test KS test
$N\!N$ 0.5210 0.5286 0.2500 0.8693
$Q$ 0.0171 0.0141 0.0155 0.0217
$C_{V}$ 0.5148 0.5577 0.2500 0.9977
$C\!E$ 0.0207 0.0156 0.0183 0.0392
$R_{nr}$ 0.0207 0.0156 0.0183 0.0392
$LES$ 0.6390 0.6575 0.2500 0.5487
Table 1: $p$-values for 1000 float points, $\alpha=0.05$

### v2

@@ 2 Results
We have demonstrated differences in the sampling itself, as well as the potential of QRNGs to yield performance advantages in terms of improved accuracy or reduced sample size in the examplary MC simulations. Higher accuracy, or reductions in sample size without loss of accuracy, should benefit every stochastic simulation (e.g., Black-Scholes, stochastic molecular dynamics), but we expect greatest benefits for problems where random numbers are essential, but do not dominate the overall computation time.
Additionally, given the observations of an increased homogeneity of samples, we anticipate that non-linear problems will enhance the observable differences between the random number generators. Further, we expect the differences to be emphasized in problems entailing estimation of rare event probabilities and general sampling of heavy-tailed distributions.
@@ 3 Conclusions
Likewise, using Buffon’s needle experiment, we demonstrated that the QRNG results pass the t-test (hypothesis: mean = $\pi$, data normally distributed) for sample sizes up to 4.54×4.54$\times$4.54 × larger than those achieved by the parallel PRNG, thus resulting in a 2$\sim$ 2$\times$ 2 × better approximation of $\pi$. If the limit is relaxed to $1/\sqrt{N}$, QRNG permits up to 3.16×3.16$\times$3.16 × larger samples, resulting in an up to 2.9×2.9$\times$2.9 × better approximation. In the process, we have demonstrated the capability of this method to distinguish the effects of points-per-sample $M$ and sample number $N$, which is not commonly found in the literature. We have also shown that the use of a QRNG results in better approximations (lower approximation errors) for sub-optimal parameter choices ($N\gg M$), enabling large reductions (8$\sim$ 8$\times$ 8 ×) of the number of samples without loss of accuracy. Finally, we demonstrated that comparisons of the effects of QRNGs and PRNGs on MC simulations should be performed with industry-standard PRNGs (e.g. Mersenne Twister or recommended pPRNGs). Additionally, based on a uniformity analysis, we assessed differences in the random sampling underpinning the MC simulations: our findings suggest that the QRNG, especially at small sample sizes, offers a better dispersion of samples than the PRNG indicating a tendency of the QRNG towards more uniformly distributed samples, closely matching theoretical expectation.
Moving beyond the investigated use cases with known solutions, we envisage, and encourage, further testing and applications of QRNGs in stochastic modelling, ranging from Bayesian inference, stochastic differential equations, optimisation, and Monte Carlo simulations to leverage the revealed advantages in all affected fields. We expect the differences between PRNGs and QRNGs to be more pronounced for non-linear problems, such as non-linear stochastic differential equations, and especially for workloads that rely upon probability estimates rather than the leading moments of a distribution. We plan to pursue a wider comparison of QRNGs, classical true random number generators (TRNGs) and a variety of PRNGs in the aforementioned fields using the methodology utilised in this paper.
@@ B.1 On the number $\pi$ approximations
However, the approximation error of method A displays no clear dependence on $M$ and is hence assumed independent of $M$.
In contrast, Buffon’s method yields larger approximation errors for $M\leq 5\cdot 10^{5}$ but displays the expected $\frac{1}{\sqrt{M}}$ convergence, resulting in a better approximation at $M=10^{6}$ pps and beyond. Hence, method B allows for a clear distinction of the effects of the sample size $N$ and the points per sample $M$, whereas for method A such a distinction is not easily made (fig. 9).